An airplane is flying with a velocity of 240m/s at an angle of 30.0 degrees. with the horizontal. when the altitude of the plane is 2.4km, a flare is released from the plane. the flare hits the target on the ground. What’s the angle?
1
Expert's answer
2019-10-17T11:11:29-0400
"h=-v\sin{30}t+0.5gt^2"

"2400=240\sin{30}t-0.5(9.8)t^2"

"t=37.5\ s"

"u_x=v\cos{30}=240\cos{30}=207.8\frac{m}{s}"

"u_y=gt-v\sin{30}=(9.8)(37.5)-240\sin{30}=247.5\frac{m}{s}"

The angle is


"\theta=\arctan{\frac{247.5}{207.8}}=50.0\degree"

below the horizontal.